### v1

# Zero-shot Class Unlearning via Layer-wise Relevance Analysis and Neuronal Path Perturbation

Abstract
In the rapid advancement of artificial intelligence, privacy protection has become crucial, giving rise to machine unlearning. Machine unlearning is a technique that removes specific data influences from trained models without the need for extensive retraining. However, it faces several key challenges, including accurately implementing unlearning, ensuring privacy protection during the unlearning process, and achieving effective unlearning without significantly compromising model performance. This paper presents a novel approach to machine unlearning by employing Layer-wise Relevance Analysis and Neuronal Path Perturbation. We address three primary challenges: the lack of detailed unlearning principles, privacy guarantees in zero-shot unlearning scenario, and the balance between unlearning effectiveness and model utility. Our method balances machine unlearning performance and model utility by identifying and perturbing highly relevant neurons, thereby achieving effective unlearning. By using data not present in the original training set during the unlearning process, we satisfy the zero-shot unlearning scenario and ensure robust privacy protection. Experimental results demonstrate that our approach effectively removes targeted data from the target unlearning model while maintaining the model’s utility, offering a practical solution for privacy-preserving machine learning.
Index Terms:
Machine unlearning, privacy preservation, model security.
I Introduction
In recent years, with the growing public attention on AI models, research has increasingly focused on privacy concerns stemming from user-model interactions. Once a model is trained, users may wish to withdraw their data and remove its influence from the model. Traditionally, this can be achieved by retraining the model, but retraining is time-consuming and inefficient for large datasets, such as those used in large language models (LLMs) [1]. The goal of unlearning is to address this issue by selectively removing the model’s memory of specific data without retraining the entire model. This can be done through data reorganization or model manipulation, allowing compliance with data deletion requests and safeguarding user privacy [2].
Machine unlearning research is primarily divided into class unlearning and sample unlearning [2, 3]. Class unlearning aims to eliminate all knowledge the model has acquired from a specific data class. This type of unlearning is often necessary when data usage policies require the deletion of certain users’ data, when the data has lost its relevance, or when it has introduced bias. In contrast, sample unlearning focuses on removing the influence of specific data samples from the model. This approach is commonly applied when users request the deletion of their data or when problematic data samples need to be excluded. Class unlearning is better suited for deleting data related to broad groups, while sample unlearning is better for individual user needs. By exploring practical ways to implement class and sample unlearning, researchers can develop more flexible privacy protection mechanisms that meet user demands without incurring unnecessary performance losses.
So far, class unlearning is particularly important and requires more research because it tackles the challenge of removing knowledge from entire categories of data, which is crucial for mitigating risks such as data misuse and unintended bias. Class unlearning methods have been applied to a wide range of AI tasks. For example, Chen et al. [4] achieved unlearning in image classification models by manipulating the model’s decision space. Ma et al. [5] implemented unlearning using neuron masks derived from gradients. Furthermore, Sun et al. [6] employed unlearning methods to make StyleGAN forget specific knowledge classes. Additionally, Chen et al. [7] proposed an unlearning framework for Graph Neural Networks (GNNs). These examples illustrate that class unlearning goes beyond simply removing individual samples—it ensures that entire problematic or sensitive data categories can be eliminated, which is essential for building fair and privacy-preserving AI systems.
Class unlearning can be approached using data reorganization and model manipulation [2]. However, both methods face a critical problem in balancing effective unlearning with maintaining model utility. The instability of the unlearning process presents challenges in two primary forms: over-unlearning and under-unlearning. Over-unlearning occurs when excessive data is removed, diminishing the model’s utility. At the same time, under-unlearning happens when insufficient data is erased, preserving the model’s usefulness but failing to achieve the unlearning objective.
As we explore the common issue of class unlearning in neural networks, a fundamental question arises: Can we achieve knowledge unlearning without compromising model utility? Delving into the challenges associated with balancing unlearning and model utility reveals three specific challenges:
  • Lack of explainable machine unlearning methods. A prominent challenge in machine unlearning is the need for more basic explanations regarding the unlearning principle. Even though users submit unlearning requests to model owners, there’s still a significant lack of clear explanations on how the unlearning methods influence the models’ parameters. This lack of clarity about unlearning principles makes it hard for users to understand how unlearning works, preventing them from trusting these methods.
  • Model owners must access users’ original data to use unlearning methods. Another critical challenge confronting machine unlearning is more privacy guarantees. Many existing methods use the users’ original model for unlearning, either by fine-tuning or by directly modifying the parameters. Using original data raises serious privacy concerns because it may expose sensitive information during model updates. Focusing on data-driven methods worsens the privacy problem, highlighting the urgent need for zero-shot machine unlearning methods.
  • The time and computational resources needed to perform machine unlearning are still excessively high. Many existing unlearning methods necessitate extensive calculations, which can involve retraining or fine-tuning models based on large datasets. This high resource demand increases operational costs and introduces delays, making it less feasible for real-world applications where efficiency and responsiveness are critical. As a result, achieving practical implementation remains a considerable hurdle.
We address the three challenges mentioned above using different approaches, thereby designing an efficient unlearning method. We investigate the relevance between the model’s neurons and the unlearning information for the first challenge. ”Relevance” refers to the contribution or importance of each neuron in the neural network to the final prediction. We identify the path of neurons with high relevance, which we define as the classification path. Our goal is to determine the neuron classification path by assessing the relevance of each neuron to the unlearning information. Through layer-wise relevance analysis, we can clearly describe the internal principles of machine unlearning.
For the second challenge, we use data not present in the training dataset to perform actions in the unlearning process, such as queries. This kind of data can be obtained in two ways. First, we can directly extract the data from the test set. When the model training is complete, the model owner will conduct performance evaluations using data that was not used for training, which can also be used for machine unlearning. Additionally, data generated by generative models can be utilized in the unlearning process. We can instruct generative models to generate data related to unlearning information and use queries or other methods to determine the relevance of neurons in the model. Through these two methods, we can ensure the privacy of users’ data during the machine unlearning process.
One last point is that our method does not require modifying the overall model parameters after the model processes the data, significantly saving time and computational resources. We analyze the data related to the unlearning information and identify the neurons that show high relevance. Due to the redundancy in neural networks, we can achieve effective unlearning by perturbing only a small subset of neurons, which minimizes the need for time and computational resources. This approach enhances the feasibility of real-world applications, enabling efficient unlearning while maintaining model utility.
In summary, we introduce layer-wise relevance analysis and neuronal path perturbation to fix the privacy and effectiveness challenges. To ensure robust privacy protection, we minimize the exposure of model owners to users’ data. Also, we implement unlearning techniques within the neural network architecture to balance unlearning effectiveness and model utility. The key contributions of this paper can be summarized as follows:
  • We propose an explainable unlearning method significantly improving unlearning speed and reducing resource consumption. By employing layer-wise relevance analysis and neuronal path perturbation, we can adjust the neuron weights highly related to unlearning information within the model. Thereby, we can achieve effective unlearning without compromising the model’s utility.
  • Our method can achieve class unlearning without accessing the user’s original data. We extract data closely related to the unlearning information from AI-generated contexts or datasets without samples from the training set to identify neurons within the neural network that are relevant to unlearning information without users’ data.
  • We demonstrate that unlearning in neural networks can be approached from the perspective of network inference by conducting experiments and analyses focused on model parameters and redundancy. Our method addresses traditional parameter adjustments and delves into the redundant characteristics within neural networks. This inference-based perspective offers new insights and methodologies for future research.
II Related Works
II-A Machine Unlearning
As the importance of data privacy protection continues to rise, class unlearning has emerged as a critical technique in machine unlearning, gaining increasing attention from researchers. However, achieving this goal is not straightforward; thoroughly exploring existing unlearning methods can provide valuable insights for our research and enhance our understanding of the challenges.
First, Bourtoule et al. [8] developed an unlearning algorithm where users find it difficult to revoke access to their data and request its deletion after sharing. They introduced the SISA training framework, which speeds up the unlearning process by strategically limiting the impact of data points during training. Chen et al. [4] proposed Boundary Unlearning, which rapidly unlearns an entire class from a DNN model by shifting its decision boundary rather than altering the model parameters. Jia et al. [9] introduced an approach to machine unlearning by shifting focus from data-centric methods to a model-based perspective, specifically using model sparsification through weight pruning. This ”prune first, then unlearn” paradigm reduces the gap between exact and approximate unlearning. Chen et al. [10] proposed a fast model debiasing framework (FMD) that identifies and removes biases in trained models using counterfactual concepts and influence functions. They integrate a machine unlearning-based strategy to efficiently debias models with minimal data and parameter updates. Kurmanji et al. [11] introduced SCRUB, a novel unlearning algorithm that excels across different unlearning applications, including removing biases (RB), resolving confusion (RC) and protecting user privacy (UP). Foster et al. [12] proposed Selective Synaptic Dampening (SSD), a novel retrain-free machine unlearning method that uses the Fisher information matrix to identify and dampen parameters critical to the data being forgotten. SSD does not require long-term training data storage and achieves performance comparable to retrain-based methods. Liu et al. [13] explored the security risks of machine unlearning by introducing backdoor attacks exploiting unlearning requests. They propose two attack methods: one where the attacker submits unlearning requests without poisoning any training data and another where the attacker poisons a few data instances with a pre-defined trigger and then activates the attack through malicious unlearning.
Moreover, Chundawat et al [14] addressed the challenge of achieving machine unlearning in scenarios without original training data, introducing the concept of zero-shot machine unlearning. They propose two methods, error minimizing-maximizing noise and gated knowledge transfer, to remove the influence of the forgotten data while preserving model performance on retained data. Tarun et al. [15] introduced a fast and scalable machine unlearning framework that uses error-maximizing noise and weight manipulation to unlearn specific data classes without accessing the entire training dataset. Their method applies an impair-repair approach, where noise is used to sharply unlearn a target class, followed by a repair step to recover overall model performance. Existing research methods have also extended to feature unlearning. Xu et al. [16] introduced a ”feature unlearning” scheme to address the limitations of existing machine unlearning methods, which typically focus on removing entire instances or classes. Recognizing the need for feature-level unlearning to maintain model utility They further proposed an approach [17] constructed an essential graph to describe relationships between critical parameters, identify parameters important to both the forgotten and retained targets and use a pruning-based technique to remove information about the target to be unlearned effectively.
Compared to the abovementioned methods, our approach focuses more on the correlation between neurons in the neural network and the unlearning data. Additionally, our method considers user privacy, as it does not require using the user’s original model, which has not received much attention in previous methods. Ultimately, our approach efficiently accomplishes the unlearning task while maintaining the model’s utility.
II-B Layer-wise Relevance Propagation
Layer-wise Relevance Propagation (LRP) aims to provide interpretability for deep learning models by attributing model predictions back to input features. It redistributes the output relevance scores to highlight which features are most influential in driving the model’s decisions, enhancing transparency and trust in its behavior.
Bach et al. [18] contributed to the development of layer-wise relevance propagation, a methodology that enables pixel-wise decomposition of classification decisions, enhancing the interpretability of automated image classification systems. Binder et al. [19] proposed an approach to extend layer-wise relevance propagation to neural networks with local renormalization layers, which is a very common product-type non-linearity in convolutional neural networks.
Today, research on LRP still holds a significant influence. Li et al. [20] proposed the Weight-dependent Baseline Layer-wise Relevance Propagation algorithm, which improves instance-level explanations by addressing limitations in existing LRP methods that disregard model weights or sample features. Bassi et al. [21] proposed ISNet, which optimizes LRP heatmaps to reduce background bias and mitigate shortcut learning in deep classifiers without increasing computational cost. Achtibat et al. [22] extended Layer-wise Relevance Propagation to attention layers in transformer models, providing a computationally efficient method for faithfully attributing input and latent representations. Hatefi et al. [23] proposed optimizing the hyperparameters of attribution methods from eXplainable AI for network pruning, demonstrating improved model compression rates for both transformer-based and convolutional networks.
Based on the above work, Layer-wise Relevance Propagation can quantitatively analyze the relevance between neurons in a neural network and specific data categories. This helps us identify which neurons are associated with the categories of unlearning data, allowing for more precise and efficient completion of the unlearning process.
III Backgrounds
TABLE I: Abbreviations and Acronyms. This table provides explanations for the abbreviations and acronyms used in the article.
$F$ Model’s prediction function
$L$ Accuracy loss function
$\theta$ Model’s parameters
$\theta^{\prime}$ Approximate parameters
$D/D^{\prime}$ Training/Updated dataset
$N_{c}$ Number of classes
$C_{i}$ Class $i$
$(x,y)$ Input-output pairs
$\eta$ Learning rate
$\nabla$ Gradient
$r$ Propagation rules
$w$ weight
$p$ neuron position
$R$ Relevance score
$l_{i}$ Layer $i$
$m_{i}$ Amount of neurons in Layer i
$m_{p}$ Amount of perturbed neurons in Layer i
$G$ Gigabyte
III-A Machine Unlearning
Machine unlearning is the process of removing specific data’s influence from a trained machine-learning model.
Consider a machine learning model with a prediction function $F$, trained on a dataset $D$ via optimizing an accuracy loss function $L$. The model’s parameters after training are represented by $\theta$. When we need to unlearn specific data $(x_{u},y_{u})$, we update the training dataset to $D^{\prime}$ by removing the data that needs to be forgotten as Equation 1.
$D^{\prime}=D\setminus\{(x_{u},y_{u})\}$ (1)
As shown in Equation 2, the updated dataset $D^{\prime}$ is then used to retrain the model, resulting in new parameters $\theta^{\prime}$ that ideally do not retain any influence from the removed data.
$\theta^{\prime}\approx\operatorname{argmin}_{\theta}\sum_{(x,y)\in D^{\prime}}L(F(x;\theta),y)$ (2)
III-B Layer-wise Relevance Propagation
Layer-wise Relevance Propagation (LRP) is a technique used to interpret the decisions made by deep neural networks. Let $F(x)$ represent the model’s prediction function for an input $x$. Given a neural network, LRP aims to attribute the prediction $F(x)$ to the individual input features by propagating relevance scores through the network layers.
For a given class $C_{i}$, the relevance score $R^{o}$ at the output layer $l_{o}$ is initialized as the prediction for that class. If the neural network outputs a probability distribution over classes, the relevance score $R_{j}^{o}$ for each output neuron $j$ corresponding to class $C_{i}$ can be initialized as:
$R_{j}^{o}=\delta_{ji}\cdot F_{C_{i}}(x),\ \delta_{ij}=\begin{cases}1&\text{if }j=i\\ 0&\text{otherwise}\end{cases}$ (3)
In Equation 3, $\delta$ is the Kronecker delta function, which is $1$ if $j=i$ and 00 otherwise, and $F_{C_{i}}(x)$ is the output of the network for class $C_{i}$. This initialization means the relevance score is set to the model’s output for the target class and zero for other classes.
The relevance scores are propagated from the output layer back to the input layer using a set of propagation rules. Let $R_{k}^{(l_{i})}$ be the relevance score of neuron $k$ in layer $l_{i}$, and we denote all the different neurons in layer $l_{i}$ as $k^{\prime}$. The relevance score is propagated to the layer $l_{i}$ from layer $l_{i+1}$ according to the following general rule:
$R_{k}^{i}=\sum_{j}\frac{z_{kj}}{\sum_{k^{\prime}}z_{k^{\prime}j}}R_{j}^{i+1},\ k\in l_{i},\ k^{\prime}\in l_{i},\ j\in l_{i+1}$ (4)
where $z_{kj}=a_{k}\cdot w_{kj}$ represents the contribution of neuron $k$ to $j$, with $a_{k}$ being the activation of neuron $k$ and $w_{kj}$ being the weight connecting neuron $k$ to $j$.
IV Methodology
IV-A Threat Model
Purpose: Our method aims to achieve class unlearning efficiently. We strive to minimize disruption to the model’s utility during unlearning. Overall, our method seeks to balance the unlearning results with the utility of the unlearned model.
Users: Users can know that their data is used to train the model and might be aware of privacy policies and data usage agreements. They can request the unlearning of their data and withdraw consent for its use in the model. After requesting machine unlearning, users may receive confirmation that their data has been removed.
Model owner: Model owners know the details of the training dataset, including data provided by each user, and understand the model’s performance metrics and how unlearning affects it. They can locate and delete specific user data upon request, retrain or adjust the model to maintain its performance and ensure compliance with relevant laws. They may need to provide users with proof of unlearning.
IV-B Overview of Our Method
Figure 1: The overview of the unlearning method. The flowchart outlines the process from data preparation and layer-wise relevance analysis to neuronal path perturbation. Specific steps include preparing data, using layer-wise relevance propagation to analyze neuron occurrence frequency, confirming the target number of neurons, and conducting neuronal path perturbation to create an unlearned model.
Our method includes several steps to achieve machine unlearning. In Figure 1, after users send unlearning requests, the model owner will prepare dataset $D^{\prime}$ for the following actions. The two primary steps are layer-wise relevance analysis and neuronal path perturbation.
Layer-wise Relevance Analysis obtain and analyze all the neurons’ relevance to the classification via LRP. According to LRP rules $r$, the predictions’ attributions to the model’s inputs are obtained through backward propagation among neurons. Therefore, we can determine the relevance of neurons to classification by comparing the relevance scores $R$ of neurons in each layer. As shown in Figure 1, it is necessary to identify the neurons’ occurrence frequency after analyzing each sample due to the unique characteristics of the samples.
Neuronal Path Perturbation perturbing the neurons in the intersection of relevant neurons to achieve unlearning. We can conduct perturbation just for the neurons in the fully connected layer before the output layer. If only the output layer is perturbed, it is similar to abandoning the model’s output of a specific class. Similarly, in Figure 1, we first confirm the target neurons quantity $m_{p}$. Then, we can perturb the target neurons to cut off the model’s neuronal classification path to achieve machine unlearning.
After these two steps, we can get the class-unlearned model without additional training. We will describe our method in the following section.
IV-C Layer-wise Relevance Analysis
A neuron network can be seen as a complex function that can finish a specific task. The networks’ predictions are the results of neurons’ internal interactions. We need to understand the internal principle of neuron networks to achieve unlearning without additional training. However, obtaining the relevance of neurons to the model’s inputs is difficult due to neuron networks’ black-box characteristics.
To fix this problem, we first apply layer-wise relevance propagation (LRP) to obtain each neuron’s internal relevance score $R$. We prepare a dataset $D^{\prime}$ that contains no users’ data to conduct LRP. The relevance $R$ is conserved through the layers, ensuring that the total relevance is preserved from the output layer back to the input layer. This process can be demonstrate in Equation 5 and 6:
$R_{j}^{l+1}=\sum_{i}R_{i\leftarrow j}^{l,l+1}$ (5)
Here, based on LRP, $R_{j}^{l+1}$ is the relevance at layer $l+1$ for neuron $j$, and $R_{i\leftarrow j}^{l,l+1}$ is the relevance propagated from neuron $j$ at layer $l+1$ to neuron $i$ at layer $l$. For layer $l$, the relevance of neuron $i$ can be computed by summing all the relevance from neurons at layer $l+1$.
$R_{i}^{l}=\sum_{j\in(l+1)}R_{i\leftarrow j}^{l,l+1}$ (6)
Via the method above, we can calculate the relevance of each neuron about input $x$ in the neural network. For the FC layer $l_{i}$, we pair the neuron positions in this layer with their relevance and sort them according to the order of the neuron positions. Assuming there are $m$ neurons in layer $l_{i}$ and the position of neuron $j$ is $p_{j}^{i}$ we can demonstrate this step as:
$List(l_{i},x)=Sort\{(R_{0}^{i},p_{0}^{i}),...,(R_{j}^{i},p_{j}^{i}),...,(R_{m-1}^{i},p_{m-1}^{i})\},$ (7)
$p_{0}^{i}<p_{1}^{i}<\ldots<p_{m-1}^{i}$
We define the top $k$ neurons with high $R$s in $List(l_{i},x)$ as high-relevance neurons. By extracting these top $k$ neurons in $List(l_{i},x)$, we obtain $List(l_{i},x,k)$ as Equation 8. Unlike Equation 7, the $(R,p)$ pairs in $List(l_{i},x,k)$ are arranged in descending order of $R$, rather than by the position of the neurons in $l_{i}$.
$List(l_{i},x,k)=\{(R_{0}^{i},p_{0}^{i}),(R_{1}^{i},p_{1}^{i}),\ldots,(R_{k-1}^{i},p_{k-1}^{i})\},$ (8)
$R_{0}^{i}\geq R_{1}^{i}\geq\ldots\geq R_{k-1}^{i},k\in\left[1,m\right]$
However, the action above only conducts relevance analysis for a single sample. We need to analyze more data to identify a set of neurons generally relevant to the data within an unlearning dataset. We use dataset $D^{\prime}$ for our analysis, which includes data from the unlearning class. These data are extracted from open-source datasets or generated by generative models to protect user privacy. We define that dataset $D^{\prime}$ contains $n$ data samples. After conducting relevance analysis in layer $l_{i}$ on each sample, we calculate the occurrence frequency $C$ of neurons to eliminate the influence of data-specific characteristics.
$C\left(p^{i}\right)=\sum_{j\in n}{Count}\left(List\left(l_{i},x_{j},k\right)\right),x_{j}\in D^{\prime},k\in\left[1,m\right]$ (9)
As shown in Equation 9, we define $C(p)$ as the occurrence frequency of neuron $p^{i}$ in all the $List(l_{i},x,k)$, which is calculated using the counting function $Count$.
For all neurons with high $R$, we obtain their occurrence frequency $C$ using the method mentioned above. Next, as in Equation 10, we sort these $C$ values in descending order to identify the strongly relevant neurons $S^{i}$ corresponding to unlearned class in layer $l_{i}$.
It is worth mentioning that due to the specificity of the images, the neurons associated with each data sample will vary. Therefore, if we denote the number of neurons that appear in all the lists as $z$, the value of z will range between $k$ and $m$.
$S^{i}=Sort\left\{C\left(p_{0}^{i}\right),C\left(p_{1}^{i}\right),...C\left(p_{z-1}^{i}\right)\right\},$ (10)
$p_{0}^{i}<p_{1}^{i}<\ldots<p_{z-1}^{i},z\in\left[k,m\right]$
IV-D Neuronal Path Perturbation
In neural networks for image classification, the initial layers typically extract low-level features such as edges and textures. The middle layers capture higher-level features like shapes and parts. In contrast, the FC layers converge the features into more abstract representations for classification so that we will perturb the FC layer $l_{i}$ for cutting off the neuronal classification path.
First, we extract the neurons’ weights that need to be perturbed. In Equation 11, we define the set of weights in layer $l_{i}$ that need to be perturbed as $W^{i}$. Formally, let $w^{i}$ represent the weights connecting layer $l_{i-1}$ to $l_{i}$. Each weight $w^{i}$ in $W^{i}$ corresponds with the position $p$ of neurons in $S^{i}$.
$W^{i}=\left\{w_{0}^{i},w_{1}^{i},...,w_{z-1}^{i}\right\},z\in\left[k,m\right]$ (11)
To facilitate the unlearning of a specific class from the neural network, we apply dropout to the neurons in layer $l_{i}$ responsible for classifying this class. As Equation 10 shows, $S^{i}$ denotes the set of neurons part of the neuronal classification path for the unlearning class in layer $l_{i}$. We implement dropout by setting the weights $w^{i}$, connecting the neurons in $S^{i}$ to the previous layer to 00 as Equation 12, effectively disrupting the classification path.
$w_{j}^{i}=0\quad\text{for all}\quad(w_{j}^{i}\in S^{i}\quad\text{and}\quad j\in\left[k,m\right])$ (12)
The neuronal path perturbation modifies the forward propagation with the dropout mask $M^{i}$ in $l_{i}$, where $M^{i}$ has entries of 0 for the weights corresponding to neurons in $S^{i}$ and $1$ otherwise, leading to the updated prediction function:
$F^{\prime}\left(x,\theta\right)=F\left(x,\theta\right)\cdot M^{i}$ (13)
By applying this targeted dropout, the network unlearns the specific class, as the classification path through layer $l_{i}$ is effectively cut off.
IV-E Relationship Between Neuron Relevance and Unlearning
In deep learning models, neurons progressively extract and represent high-dimensional features of the input data through multi-layer structures and nonlinear activation functions. Neurons’ relevance to specific data samples reflects their role in processing these samples, indicating how well they capture the essential features of these samples. The relevance of neurons to data samples directly impacts the model’s classification performance on these samples.
Different neurons may have varying sensitivities to various features. Some neurons are highly sensitive to specific features that are key distinguishing characteristics of the data samples’ categories. Therefore, the activation states of these neurons significantly determine the model’s classification decisions. This implies that the model’s classification relies on neurons that can effectively extract and represent the input data’s features.
Identifying neurons with high relevance to the unlearning data becomes crucial in machine unlearning. By targeting and interfering with these specific neurons, we can effectively disrupt their ability to capture and represent the features of the unlearning data. This disruption can be achieved by altering the weights or applying noise to the outputs of these neurons. Consequently, the model’s ability to classify the unlearning data will be impaired, achieving the goal of unlearning.
This targeted interference should be designed to affect the neurons’ role in processing minimally and classifying the remaining data. Because the neurons with high relevance to the unlearning data are specifically identified and targeted, the model’s performance on the remaining data can be preserved. This selective approach ensures that the model maintains its accuracy and robustness for the data not subject to unlearning, thereby achieving effective unlearning without compromising the overall model performance.
V Expriment
V-A Datasets and Setups
In our experiments, we utilized various datasets and models. We will introduce our work separately for image classification tasks.
MNIST Dataset [24]. The MNIST dataset is a large database of handwritten digits commonly used for training various image processing systems. The dataset contains $60,000$ training images and $10,000$ testing images, each with a resolution of $28$x28282828 pixels.
CIFAR-10 Dataset [25]. The CIFAR-10 dataset consists of $60,000$ color images in $10$ different classes, with $6,000$ images per class. Each image has a resolution of $32$x32323232 pixels and is widely used for training and evaluating machine learning algorithms in image classification.
CIFAR-100 Dataset [25]. The CIFAR-100 dataset is similar to CIFAR-10 but contains $100$ classes, each with $600$ images, making a total of $60,000$ images. The images are also $32$x32323232 pixels in size, and this dataset is often used for more challenging tasks in image recognition, requiring algorithms to learn finer distinctions among categories.
mini-ImageNet Dataset [26]. mini-ImageNet is a massive database of images that are hand-annotated and categorized into $1,000$ different classes, ranging from objects to scenes to abstract concepts. It’s like an enormous visual dictionary for computers, particularly in image recognition and classification.
Our experiments preprocessed the CIFAR-10, CIFAR-100, and mini-ImageNet datasets to adjust their resolutions, making them more aligned with real-world image scenarios. Table II provides each dataset’s image resolution and the corresponding classification model. Furthermore, to verify that our method can address Challenge $2$ by enabling the model owner to unlearn the unlearning class without accessing the training data, we used data from the test set of each dataset as samples for layer-wise relevance analysis.
TABLE II: Datasets with Image Resolution and Classification Model
Dataset Image Resolution Classification Model
MNIST 28x28 ALLCNN
CIFAR-10 224x224 ResNet-50
CIFAR-100 224x224 ResNet-50
mini-ImageNet 256x256 VGG-16
V-B Evaluation Metrics
V-B1 Evaluation Metrics for Images unlearning
In our experiments on image unlearning, we utilized a range of intuitive and precise evaluation metrics to illustrate the effectiveness of our proposed approach. The metrics employed are outlined below:
Unlearning class accuracy ($A_{t}$). The accuracy of the model on the unlearning class, $A_{t}$, reflects the model’s performance during the forgetting process. A high $A_{t}$ suggests that the model struggles to remove the knowledge of the target class from its memory. Conversely, a low $A_{t}$ indicates successful unlearning to the target class.
Global accuracy ($A_{g}$). The model’s average accuracy on the remaining classes is denoted as $A_{g}$. $A_{g}$ can be obtained by calculating the model’s accuracy on the dataset after removing the data from the target unlearned class. A high $A_{g}$ reflects the improved utility of the unlearned model for the remaining tasks. Furthermore, a lower $A_{g}$ signifies a decrease in the model’s utility after unlearning.
Forgetting rate ($Fr$). The forgetting rate of the unlearned model is demonstrated as $Fr$. $Fr$ is calculated by the proportion of ”data not trained by the model” in the results of the membership inference attack. A higher $Fr$ indicates that more unlearning data has been forgotten.
Gigabyte ($G$). In this context, Gigabyte (GB) is employed to measure the computational resource consumption of GPU memory. A Gigabyte represents $10^{9}$ bytes of data. We denote Gigabyte as $G$. Higher $G$ signifies higher resource consumption.
Time. The time spent by the unlearning process is demonstrated as $T$. $T$ records the time taken from the model owner’s initiation of unlearning until completion. The method requires less $T$ for machine unlearning, indicating its greater efficiency.
V-B2 Comparing methods list
  1. 1.
    Retrain from Scratch involves completely retraining the model from its initial state after removing specific data points. This method ensures that the influence of the removed data is entirely eradicated from the model.
  2. 2.
    Random Labels [27] using random labels for unlearning involves randomizing the data labels of the target class, making it difficult for the model to recognize and learn the features of these data, thereby achieving effective unlearning of that class.
  3. 3.
    Amnesiac Unlearning [28] involves tracking the parameter updates from each batch during training and, upon receiving a data removal request, selectively undoing the parameter updates from the batches containing the sensitive data, thereby efficiently removing specific learned data from the neural network model.
  4. 4.
    Boundary Shrink [4] achieves machine unlearning by identifying and targeting the decision boundary of the class to be forgotten. It reassigns labels of unlearning samples to nearby incorrect classes, shrinking the decision space for the unlearning class.
  5. 5.
    Boundary Expanding [4] facilitates rapid unlearning by introducing a shadow class to disperse the activations of unlearning samples within an expanded decision space. After finetuning, the shadow class is pruned, removing the influence of unlearning data.
  6. 6.
    Neuronal Path Noise refers to introducing Gaussian or Laplacian noise to neurons identified as relevant to the unlearning dataset using the LRP method. This added noise aims to disrupt the model’s ability to classify the unlearning dataset accurately.
V-C Performance Evaluation
V-C1 Changes in quantitative metrics
Information about specific classes is relatively easy to obtain in image classification tasks. We employ the proposed unlearning method to unlearn a specific class in the MNIST, CIFAR-10, CIFAR-100, and mini-ImageNet datasets.
We choose retraining as a baseline for unlearning. The performance of the retrained model in the unlearning class provides a baseline for assessing the effectiveness of unlearning methods. Then, We will apply the $\epsilon$-rule to LRP to analyze neuron relevance to verify the unlearning method’s effectiveness. $\epsilon$-rule involves adding a small term $\epsilon$ to the denominator in the relevance propagation formula to avoid zero division and control the redistribution of relevance scores.
For the MNIST dataset, we perform unlearning on class $1$. We extract $36$ images belonging to class $1$ and have the model classify them while analyzing the relevance patterns among its neurons.
TABLE III: Class Unlearning Performance on the MNIST dataset
Method $A_{t}$ $A_{g}$ $Fr$ $G$ $Time$
Retrain from scratch $0.00$ $0.97$ $1.00$ $0.74$ $71.37s$
Random labels $0.00$ $0.97$ $1.00$ $0.70$ $15.64s$
Amnesiac $0.38$ $0.98$ $0.78$ $0.44$ $38.72s$
Boundary expanding $0.13$ $0.95$ $0.91$ $0.65$ $69.21s$
Boundary shrink $0.05$ $0.58$ $1.00$ $0.71$ $18.55s$
Neuronal path noise (GN) $0.35$ $0.68$ $0.61$ $0.64$ $0.77s$
Neuronal path noise (LN) $0.21$ $0.53$ $0.78$ $0.64$ $0.83s$
Neuronal path perturbation $0.00$ $0.97$ $1.00$ $0.64$ $0.68s$
In Table III, after applying our unlearning method, the model’s accuracy $A_{t}$ on class $1$ dropped to $0.00$, matching the $A_{t}$ observed after retraining. This demonstrates that the model has completely lost its ability to classify data from the unlearning class. Disrupting the neurons associated with the unlearning class makes it possible to cut off the neural network’s classification path.
While cutting off the model’s classification path for the unlearning class, we preserve the model’s ability to classify the remaining classes. For example, after retraining from scratch, the model achieves an overall accuracy $A_{g}$ of $0.97$ on the remaining classes. By applying our machine unlearning method, the model maintains an $A_{g}$ of $0.97$ on these classes, demonstrating strong performance. This shows that our method effectively unlearns specific class information without additional training and preserves the model’s classification capability.
Our method also excels in unlearning verification through membership inference attacks. After retraining, the new model achieves a perfect forgetting rate of $1.00$ for the unlearning class, indicating that it is entirely unaffected by the unlearning data. Our approach also achieves a forgetting rate of $1.00$, demonstrating that severing the classification path in the neural network effectively accomplishes the class unlearning task.
Also, our method performs well in terms of computational resource consumption. When using retraining to achieve unlearning, the GPU memory consumption $G$ was $0.74$, which took $71.37$ seconds. In contrast, we completed class unlearning in just $0.68$ seconds using our unlearning method. Additionally, GPU resource consumption was only $0.64$, and we used just $36$ images from class $1$ for unlearning. These results show that our unlearning method reduces resource consumption without additional training.
TABLE IV: Class Unlearning Performance on the CIFAR-10 dataset
Method $A_{t}$ $A_{g}$ $Fr$ $G$ $Time$
Retrain from scratch $0.00$ $0.96$ $1.00$ $6.89$ $484.17s$
Random labels $0.00$ $0.90$ $0.99$ $6.89$ $107.99s$
Amnesiac $0.91$ $0.84$ $0.13$ $1.75$ $3.30s$
Boundary expanding $0.83$ $0.89$ $0.01$ $4.07$ $150.22s$
Boundary shrink $0.19$ $0.58$ $0.09$ $6.76$ $20.41s$
Neuronal path noise (GN) $0.95$ $0.79$ $0.00$ $2.82$ $14.17s$
Neuronal path noise (LN) $0.99$ $0.71$ $0.00$ $2.82$ $15.51s$
Neuronal path perturbation $0.03$ $0.97$ $0.99$ $2.82$ $14.44s$
The results in Table IV demonstrate that Neuronal Path Perturbation is the most effective method for class unlearning on the CIFAR-10 dataset, achieving an excellent balance between unlearning performance and model utility. With an accuracy of $A_{t}=0.03$ on the target class, it ensures nearly complete unlearning, comparable to retraining from scratch ($A_{t}=0.00$) and random labels ($A_{t}=0.00$). Its forgetting rate ($Fr=0.99$) is almost as high as the ideal standard of $Fr=1.00$, indicating a solid unlearning effect.
At the same time, Neuronal Path Perturbation maintains high accuracy on the remaining classes ($A_{g}=0.97$), outperforming most other methods and closely matching the retrained model’s $A_{g}=0.96$. This reflects its ability to retain model utility while effectively unlearning the target class.
Moreover, Neuronal Path Perturbation is highly efficient in terms of computational resources. It consumes only $2.82$ GB of GPU memory, significantly less than retraining ($6.89$ GB), and takes just $14.44$ seconds to complete the unlearning process, a fraction of the $484.17$ seconds required for retraining. This makes it a highly practical and efficient choice for unlearning tasks.
TABLE V: Class Unlearning Performance on the CIFAR-100 dataset
Method $A_{t}$ $A_{g}$ $Fr$ $G$ $Time$
Retrain from scratch $0.00$ $0.82$ $1.00$ $6.89$ $607.89s$
Random labels $0.00$ $0.82$ $1.00$ $6.89$ $122.55s$
Amnesiac $0.03$ $0.70$ $0.29$ $1.75$ $18.45s$
Boundary expanding $0.77$ $0.81$ $0.00$ $4.07$ $83.36s$
Boundary shrink $0.14$ $0.45$ $0.07$ $7.02$ $21.20s$
Neuronal path noise (GN) $0.21$ $0.56$ $0.00$ $2.83$ $0.28s$
Neuronal path noise (LN) $0.00$ $0.29$ $0.00$ $2.83$ $0.37s$
Neuronal path perturbation $0.01$ $0.82$ $1.00$ $2.77$ $0.53s$
The results for class unlearning on the CIFAR-100 dataset highlight that our method, Neuronal Path Perturbation, offers the best balance between achieving effective unlearning and preserving model utility. With an accuracy of $A_{t}=0.01$ on the target class, this method achieves near-complete unlearning, comparable to retraining from scratch and random labels (both $A_{t}=0.00$). Its forgetting rate ($Fr=1.00$) is identical to these methods, eliminating the learned information about the target class.
Despite this solid unlearning performance, Neuronal Path Perturbation preserves a high level of accuracy on the remaining classes ($A_{g}=0.82$), matching that of the retrained model. This shows that the method effectively retains the model’s utility across the non-target classes, making it highly efficient in maintaining overall performance.
Regarding resource efficiency, Neuronal Path Perturbation consumes only $2.77$ GB of GPU memory, much lower than retraining ($6.89$ GB). It requires just $0.53$ seconds to complete the unlearning process—an impressive speed compared to the $607.89$ seconds for retraining. These results make it a highly efficient and practical solution for class unlearning tasks on CIFAR-100.
TABLE VI: Class Unlearning Performance on the mini-ImageNet dataset
Method $A_{t}$ $A_{g}$ $Fr$ $G$ $Time$
Retrain from scratch $0.00$ $0.83$ $1.00$ $6.51$ $2865.87s$
Random labels $0.17$ $0.14$ $0.74$ $2.47$ $582.32s$
Amnesiac $0.53$ $0.80$ $0.44$ $7.61$ $20.09s$
Boundary expanding $0.03$ $0.64$ $1.00$ $7.68$ $697.89s$
Boundary shrink $0.35$ $0.67$ $0.76$ $9.13$ $4.03s$
Neuronal path noise (GN) $0.91$ $0.00$ $1.00$ $0.93$ $1.16s$
Neuronal path noise (LN) $0.14$ $0.00$ $1.00$ $0.93$ $1.37s$
Neuronal path perturbation $0.00$ $0.83$ $1.00$ $0.93$ $1.35s$
For a VGG-$16$ model pre-trained on the ImageNet dataset, we employed machine unlearning using $33$ images related to the unlearning class. The mini-ImageNet dataset also has $1,000$ classes. While each class has fewer images than ImageNet, it’s still a good way to test if unlearning methods work for real-world multi-class tasks. Our experimental purpose is to make the pre-trained VGG model forget the knowledge about class 00.
As shown in Table VI, after applying our unlearning method, the model’s $A_{t}$ dropped to $0.00$, while $A_{g}$ remained at $0.83$. Our method’s unlearning results are consistent with those obtained from validating the retrained model. These results indicate that our method enabled the model to unlearn information specific to the unlearning class.
Similarly, employing a membership inference attack based on SVM for unlearning verification on VGG yielded favorable results. After retraining the model, the $Fr$ is $1.00$. Moreover, the $Fr$ of the unlearned model obtained using our method is also $1.00$, indicating that the model has wholly unlearned the knowledge of the unlearning class data. The results demonstrate that perturbing the neuronal classification path is sufficient to make the model lose its classification ability for a specific class.
When retraining the model, we set the training batch size to $32$ to achieve better results. The retraining process consumed approximately $6.51$G GPU memory. In contrast, our method only involves model inference, consuming just $0.93$G GPU memory, significantly less than the computational resources required for retraining. Moreover, from the time cost perspective, our method only requires $0.68$ seconds to complete unlearning, whereas retraining takes $2865.87$ seconds. This demonstrates that our method is highly efficient.
V-C2 Changes in attribution results
After evaluated our method with various models and datasets, we conducted attribution experiments on both the target model and the unlearned model using different data. The results show that the unlearned model exhibits more noticeable attribution changes for data in the unlearning class.
In Figure 2, the red areas represent parts of the image that positively contribute to the model’s classification, while the blue areas indicate negative contributions. Specifically, Figure 2(a) and Figure 2(b) show the changes in attribution for the model before and after unlearning the unlearning class data. In Figure 2(b), the model still extracts the main information from the image. However, there are noticeably more pixels with negative contributions (blue pixels) compared to Figure 2(a). This shows that the key information used for classification does not follow the original classification path in the unlearned model. Therefore, the model has successfully achieved unlearning.
In Figure 2(c) and Figure 2(d), we can see that after unlearning, the model’s attribution results for the same non-unlearning class data show almost no changes. Additionally, the model’s classification results for the non-unlearning class remain unchanged. This indicates that our method successfully performs unlearning while preserving the model’s utility on non-unlearning class data.
(a) The attribution of data from the unlearning class before unlearning, using various propagation rules.
(b) The attribution of data from the unlearning class after unlearning, using various propagation rules.
(c) The attribution of data from the non-unlearning class before unlearning, using various propagation rules.
(d) The attribution of data from the non-unlearning class after unlearning, using various propagation rules.
Figure 2: The attribution of the model’s classification on unlearning and non-unlearning class data from MNIST, before and after unlearning, using different propagation rules.
(a) The attribution of data from the unlearning class before unlearning, using various propagation rules.
(b) The attribution of data from the unlearning class after unlearning, using various propagation rules.
(c) The attribution of data from the non-unlearning class after unlearning, using various propagation rules.
(d) The attribution of data from the non-unlearning class after unlearning, using various propagation rules.
Figure 3: The attribution of the model’s classification on unlearning and non-unlearning class data from CIFAR-10, before and after unlearning, using different propagation rules.
In Figures 3(a) and 3(b), we compare the attributions of the ResNet50 model for unlearning class data in the CIFAR-10 dataset. Figure 3(a) shows the model’s attributions for unlearning class data before unlearning. After unlearning, Figure 3(b) displays changes in the attribution distribution. In Figure 3(b), when using attribution rules such as $epsilon$, $gamma+epsilon$, and $alpha1beta0$, we observe that more pixels with altered attributions in the unlearning data influence the classification of the unlearned model.
Meanwhile, the unlearned model’s classification on non-unlearning class data remains unaffected. In Figures 3(c) and 3(d), we observe that the attributions for non-unlearning data show no significant differences between the original model and the unlearned model. This is because the classification paths for non-unlearning class data remain undisturbed, allowing normal forward and backward propagation. These results demonstrate that the unlearned model preserves its utility on non-unlearning data.
When performing class unlearning on data from CIFAR-100, the changes in attribution maps are relatively small, but the specified categories can still be effectively forgotten. In Figures 4(a) and 4(b), we can see that the model extracts the main objects in the images both before and after unlearning, and the attribution results under the same rules have also changed. We believe that the minimal changes in attribution results are because, after training on a dataset with many classes, the model has lower parameter redundancy and only needs to perturb a few neurons to achieve unlearning. When a large number of neurons remain, the model can more comprehensively analyze the contribution of more pixels to image classification.
Similarly to the previous results, our unlearned model shows nearly no difference in attribution results for non-unlearning class data in CIFAR-10 compared to before unlearning. Our method effectively preserves the utility of the unlearned model.
(a) The attribution of data from the unlearning class before unlearning, using various propagation rules.
(b) The attribution of data from the unlearning class after unlearning, using various propagation rules.
(c) The attribution of data from the non-unlearning class after unlearning, using various propagation rules.
(d) The attribution of data from the non-unlearning class after unlearning, using various propagation rules.
Figure 4: The attribution of the model’s classification on unlearning and non-unlearning class data from CIFAR-100, before and after unlearning, using different propagation rules.
In Figure 5, we can see the attribution results of our method validated on the mini-ImageNet dataset. Figure 5(a) shows the attribution of data from the unlearning class before unlearning. After unlearning, in Figure 5(b), we can observe significant changes in pixel details. This is because neurons strongly related to the unlearning class data have been perturbed, causing the model to fail to classify normally.
Figure 5(c) and Figure 5(d) didn’t show any difference that can be directly observed. This is because the classification paths related to non-unlearning class data didn’t get perturbed during machine unlearning, indicating that our unlearning method is effective and robust.
(a) The attribution of data from the unlearning class before unlearning, using various propagation rules.
(b) The attribution of data from the unlearning class after unlearning, using various propagation rules.
(c) The attribution of data from the non-unlearning class after unlearning, using various propagation rules.
(d) The attribution of data from the non-unlearning class after unlearning, using various propagation rules.
Figure 5: The attribution of the model’s classification on unlearning and non-unlearning class data from mini-ImageNet, before and after unlearning, using different propagation rules.
V-D Parameters Evaluation
After validating our method, we attempted to use different parameters to verify the effectiveness of our approach. We will adjust the algorithm parameters from four perspectives to validate the effectiveness and universality of our unlearning algorithm. These perspectives are: different propagation rules, different amount of perturbed neurons, different amount of analyzed neurons.
V-D1 Differrnt Propagation Rules
We use different rules to analyze and perturb different numbers of neurons to reveal the effectiveness and robustness of our approach. The relevance propagation rules that we chose are as follows:
  1. 1.
    Epsilon rule is one of the basic LRP rules. It helps to prevent numerical instability by ensuring that small denominators do not result in large relevance scores.
  2. 2.
    Gamma+epsilon rule is an extension of the epsilon rule, where a positive bias $\gamma$ is added to enhance the relevance of positive contributions.
  3. 3.
    Alpha1beta0 rule explicitly separates the positive and negative contributions, using only the positive contributions for relevance propagation.
  4. 4.
    Alpha2beta1 rule is a generalized form of the alpha1beta0 rule, where both positive and negative contributions are considered but with different weights.
  5. 5.
    PatternNet $S(x)$ is designed to provide a more interpretable gradient by isolating the patterns in the input data the neural network has learned to respond to.
  6. 6.
    PatternNet $S(x)_{+-}$ is a variant of PatternNet that distinguishes between positive and negative patterns, allowing for a more detailed analysis.
  7. 7.
    PatternAttribution $S(x)$ extends PatternNet by attributing the network’s output prediction to the input features, similar to LRP but focusing on learned patterns.
  8. 8.
    PatternAttribution $S(x){+-}$ rule is similar to PatternAttribution but allows for the separation of positive and negative contributions, similar to the $S(x){+-}$ rule in PatternNet.
Table 6(a) illustrates the performance of our method across different relevance propagation rules, focusing on the accuracy of the model on remaining data ($A_{g}$) and unlearning data ($A_{t}$) as the proportion of perturbed neurons ($M_{p}$) in a single fully-connected layer increases.
For the epsilon rule, the model starts with a high accuracy of $0.99$ when no neurons are perturbed, which drops drastically to $0.001$ when $20\%$ of the neurons are perturbed. Despite this, the accuracy of the remaining data remains high at $0.97$, showcasing the robustness of the model. The gamma+epsilon rule follows a similar pattern, with the $A_{t}$ decreasing from $0.99$ to $0.001$ as the perturbation level increases to $20\%$, while the accuracy on the remaining data remains stable at $0.97$.
As shown in table 6(a), the alpha1beta0 and alpha2beta1 rules exhibit similar behavior, with the accuracy on unlearning data dropping from $0.99$ to $0.001$ as $M_{p}$ increases, and the remaining data accuracy remaining high at $0.97$. These trends confirm the effectiveness and robustness of the method across different relevance propagation rules.
In contrast, PatternNet $S(x)$ shows a more gradual decrease in $A_{t}$, from $0.99$ to $0.19$ as $M_{p}$ increases to $20\%$. The accuracy of the remaining data decreases slightly to $0.95$ when $20\%$ of neurons are perturbed, indicating a slight reduction in utility. Similarly, PatternNet $S(x){+-}$ demonstrates a moderate unlearning effect, with the $A_{t}$ dropping from $0.99$ to $0.28$ as $M_{p}$ increases, while the remaining data accuracy decreases slightly to $0.96$, still maintaining robustness.
PatternAttribution $S(x)$ maintains a relatively high $A_{t}$, decreasing only to $0.83$ when $20\%$ of neurons are perturbed, indicating less effective unlearning. However, the remaining data accuracy remains consistent at $0.97$, suggesting high robustness but limited unlearning effectiveness. On the other hand, PatternAttribution $S(x){+-}$ shows a sharp decrease in unlearning accuracy from $0.99$ to $0.00$ as the perturbation level increases, similar to the epsilon and gamma+epsilon rules. Despite this, the accuracy of the remaining data decreases only slightly to $0.95$, demonstrating strong unlearning effectiveness with reasonable robustness.
TABLE VII: Performance of the method on the MNIST dataset when Analyzing Neurons’ Relevance
$r$ $M_{p}$ $A_{g}$/$A_{t}$ $0\%$ $4\%$ $8\%$ $12\%$ $16\%$ $20\%$
epsilon $0.98$ / $0.99$ $0.98$ / $0.99$ $0.97$ / $0.64$ $0.96$ / $0.13$ $\mathbf{0.96}$ / $\mathbf{0.005}$ $\mathbf{0.97}$ / $\mathbf{0.001}$
gamma+epsilon $0.98$ / $0.99$ $0.98$ / $0.99$ $0.98$ / $0.59$ $0.96$ / $0.13$ $\mathbf{0.96}$ / $\mathbf{0.001}$ $\mathbf{0.97}$ / $\mathbf{0.001}$
alpha1beta0 $0.98$ / $0.99$ $0.98$ / $0.99$ $0.97$ / $0.56$ $0.96$ / $0.13$ $\mathbf{0.96}$ / $\mathbf{0.003}$ $\mathbf{0.97}$ / $\mathbf{0.001}$
alpha2beta1 $0.98$ / $0.99$ $0.98$ / $0.99$ $0.97$ / $0.61$ $0.96$ / $0.13$ $\mathbf{0.96}$ / $\mathbf{0.007}$ $\mathbf{0.97}$ / $\mathbf{0.001}$
PatternNet $S(x)$ $0.98$ / $0.99$ $0.98$ / $0.99$ $0.98$ / $0.95$ $0.97$ / $0.59$ $0.97$ / $0.38$ $0.95$ / $0.19$
PatternNet $S(x)_{+-}$ $0.98$ / $0.99$ $0.98$ / $0.99$ $0.98$ / $0.93$ $0.97$ / $0.62$ $0.96$ / $0.52$ $0.96$ / $0.28$
PatternAttribution $S(x)$ $0.98$ / $0.99$ $0.98$ / $0.99$ $0.97$ / $0.99$ $0.97$ / $0.97$ $0.97$ / $0.93$ $0.97$ / $0.83$
PatternAttribution $S(x){+-}$ $0.98$ / $0.99$ $0.98$ / $0.99$ $0.97$ / $0.55$ $0.96$ / $0.21$ $\mathbf{0.96}$ / $\mathbf{0.02}$ $\mathbf{0.95}$ / $\mathbf{0.00}$
(a) Top 50 Neurons’ Relevance
$r$ $M_{p}$ $A_{g}$/$A_{t}$ $0\%$ $4\%$ $8\%$ $12\%$ $16\%$ $20\%$ $24\%$
epsilon $0.98$ / $0.99$ $0.98$ / $0.99$ $0.98$ / $0.99$ $0.97$ / $0.89$ $0.97$ / $0.34$ $\mathbf{0.96}$ / $\mathbf{0.09}$ $\mathbf{0.97}$ / $\mathbf{0.07}$
gamma+epsilon $0.98$ / $0.99$ $0.98$ / $0.99$ $0.98$ / $0.99$ $0.97$ / $0.89$ $0.96$ / $0.38$ $\mathbf{0.96}$ / $\mathbf{0.09}$ $\mathbf{0.97}$ / $\mathbf{0.07}$
alpha1beta0 $0.98$ / $0.99$ $0.98$ / $0.99$ $0.98$ / $0.99$ $0.97$ / $0.89$ $0.97$ / $0.36$ $\mathbf{0.96}$ / $\mathbf{0.09}$ $\mathbf{0.96}$ / $\mathbf{0.08}$
alpha2beta1 $0.98$ / $0.99$ $0.98$ / $0.99$ $0.98$ / $0.99$ $0.97$ / $0.89$ $0.96$ / $0.40$ $\mathbf{0.96}$ / $\mathbf{0.09}$ $\mathbf{0.97}$ / $\mathbf{0.07}$
PatternNet $S(x)$ $0.98$ / $0.99$ $0.98$ / $0.99$ $0.98$ / $0.95$ $0.96$ / $0.59$ $0.96$ / $0.38$ $0.95$ / $0.19$ $\mathbf{0.95}$ / $\mathbf{0.12}$
PatternNet $S(x)_{+-}$ $0.98$ / $0.99$ $0.98$ / $0.99$ $0.98$ / $0.93$ $0.97$ / $0.62$ $0.96$ / $0.52$ $0.96$ / $0.28$ $0.95$ / $0.28$
PatternAttribution $S(x)$ $0.98$ / $0.99$ $0.98$ / $0.99$ $0.98$ / $0.99$ $0.97$ / $0.97$ $0.97$ / $0.93$ $0.97$ / $0.83$ $0.97$ / $0.75$
PatternAttribution $S(x){+-}$ $0.98$ / $0.99$ $0.98$ / $0.99$ $0.97$ / $0.55$ $0.96$ / $0.21$ $\mathbf{0.96}$ / $\mathbf{0.02}$ $\mathbf{0.95}$ / $\mathbf{0.00}$ $\mathbf{0.96}$ / $\mathbf{0.00}$
(b) Top 100 Neurons’ Relevance
We evaluated our method on the CIFAR-10 dataset using different rules in the layer-wise relevance propagation framework. Table 6(c) shows that our method can successfully achieve unlearning across all tested rules. Specifically, at $M_{p}$ values of $25\%$ and $30\%$, the four rules (epsilon, gamma+epsilon, alpha1beta0, and alpha2beta1) maintain a high accuracy of $A_{g}\approx 0.97$ on the non-target classes, indicating that the model’s performance on non-target classes remains largely unaffected. Meanwhile, the target class accuracy $A_{t}$ drops significantly, reaching values as low as $0.03$ to $0.00$ for the epsilon, gamma+epsilon, and alpha-series rules at $25\%$ and $30\%$. This demonstrates that our unlearning method effectively erases information from the target class under these conditions.
In Table 6(e), on the CIFAR-100 dataset, the results further validate the effectiveness of our method under different rules. At $M_{p}$ values of $10\%$ and $12.5\%$, all rules (epsilon, gamma+epsilon, alpha1beta0, and alpha2beta1) maintain $A_{g}=0.81$, indicating stable performance on the non-target classes. Meanwhile, the target class accuracy $A_{t}$ drops to as low as $0.01$ and $0.00$, demonstrating that our method successfully unlearns the target class information under these rules.
In the mini-ImageNet dataset, the results demonstrate that our method effectively achieves unlearning across different rules. At $M_{p}$ values of $10\%$ and $12\%$, all rules (epsilon, gamma+epsilon, alpha1beta0, and alpha2beta1) maintain $A_{g}=0.83$, indicating that the performance on non-target classes remains unaffected. Meanwhile, the target class accuracy $A_{t}$ drops to $0.00$ at $10\%$ and $12\%$, showing that our method successfully unlearns the target class information under these conditions.
TABLE VIII: Performance of the method on the CIFAR-10 dataset when Analyzing Neurons’ Relevance
$r$ $M_{p}$ $A_{g}$/$A_{t}$ $0\%$ $10\%$ $15\%$ $20\%$ $25\%$ $30\%$
epsilon $0.98$ / $0.99$ $0.98$ / $0.67$ $0.98$ / $0.34$ $0.98$ / $0.11$ $\mathbf{0.97}$ / $\mathbf{0.03}$ $\mathbf{0.97}$ / $\mathbf{0.06}$
gamma+epsilon $0.98$ / $0.99$ $0.98$ / $0.67$ $0.98$ / $0.34$ $0.98$ / $0.11$ $\mathbf{0.97}$ / $\mathbf{0.01}$ $\mathbf{0.93}$ / $\mathbf{0.00}$
alpha1beta0 $0.98$ / $0.99$ $0.98$ / $0.67$ $0.98$ / $0.34$ $0.98$ / $0.11$ $\mathbf{0.97}$ / $\mathbf{0.02}$ $\mathbf{0.93}$ / $\mathbf{0.01}$
alpha2beta1 $0.98$ / $0.99$ $0.98$ / $0.67$ $0.98$ / $0.33$ $0.98$ / $0.11$ $\mathbf{0.97}$ / $\mathbf{0.02}$ $\mathbf{0.95}$ / $\mathbf{0.01}$
(c) Top 150 Neurons’ Relevance
$r$ $M_{p}$ $A_{g}$/$A_{t}$ $0\%$ $10\%$ $15\%$ $20\%$ $25\%$ $30\%$
epsilon $0.98$ / $0.99$ $0.98$ / $0.67$ $0.98$ / $0.33$ $0.98$ / $0.11$ $\mathbf{0.97}$ / $\mathbf{0.01}$ $\mathbf{0.95}$ / $\mathbf{0.00}$
gamma+epsilon $0.98$ / $0.99$ $0.98$ / $0.67$ $0.98$ / $0.33$ $0.97$ / $0.11$ $\mathbf{0.97}$ / $\mathbf{0.01}$ $\mathbf{0.95}$ / $\mathbf{0.00}$
alpha1beta0 $0.98$ / $0.99$ $0.98$ / $0.67$ $0.98$ / $0.33$ $0.98$ / $0.11$ $\mathbf{0.97}$ / $\mathbf{0.02}$ $\mathbf{0.96}$ / $\mathbf{0.00}$
alpha2beta1 $0.98$ / $0.99$ $0.98$ / $0.67$ $0.98$ / $0.33$ $0.98$ / $0.11$ $\mathbf{0.97}$ / $\mathbf{0.02}$ $\mathbf{0.95}$ / $\mathbf{0.00}$
(d) Top 200 Neurons’ Relevance
TABLE IX: Performance of the method on the CIFAR-100 dataset when Analyzing Neurons’ Relevance
$r$ $M_{p}$ $A_{g}$/$A_{t}$ $0\%$ $7.5\%$ $10\%$ $12.5\%$
epsilon $0.82$ / $0.96$ $0.81$ / $0.07$ $\mathbf{0.81}$ / $\mathbf{0.01}$ $\mathbf{0.81}$ / $\mathbf{0.00}$
gamma+epsilon $0.82$ / $0.96$ $0.81$ / $0.07$ $\mathbf{0.81}$ / $\mathbf{0.01}$ $\mathbf{0.81}$ / $\mathbf{0.00}$
alpha1beta0 $0.82$ / $0.96$ $0.81$ / $0.05$ $\mathbf{0.81}$ / $\mathbf{0.01}$ $\mathbf{0.81}$ / $\mathbf{0.00}$
alpha2beta1 $0.82$ / $0.96$ $0.81$ / $0.07$ $\mathbf{0.81}$ / $\mathbf{0.01}$ $\mathbf{0.81}$ / $\mathbf{0.00}$
(e) Top 150 Neurons’ Relevance
$r$ $M_{p}$ $A_{g}$/$A_{t}$ $0\%$ $7.5\%$ $10\%$ $12.5\%$
epsilon $0.82$ / $0.96$ $0.81$ / $0.08$ $\mathbf{0.81}$ / $\mathbf{0.01}$ $\mathbf{0.81}$ / $\mathbf{0.00}$
gamma+epsilon $0.82$ / $0.96$ $0.81$ / $0.10$ $\mathbf{0.81}$ / $\mathbf{0.01}$ $\mathbf{0.81}$ / $\mathbf{0.00}$
alpha1beta0 $0.82$ / $0.96$ $0.81$ / $0.09$ $\mathbf{0.81}$ / $\mathbf{0.01}$ $\mathbf{0.81}$ / $\mathbf{0.00}$
alpha2beta1 $0.82$ / $0.96$ $0.81$ / $0.09$ $\mathbf{0.81}$ / $\mathbf{0.01}$ $\mathbf{0.81}$ / $\mathbf{0.00}$
(f) Top 200 Neurons’ Relevance
TABLE X: Performance of the method on the mini-ImageNet dataset when Analyzing Neurons’ Relevance
$r$ $M_{p}$ $A_{g}$/$A_{t}$ $0\%$ $2\%$ $4\%$ $8\%$ $10\%$ $12\%$
epsilon $0.83$ / $0.88$ $0.83$ / $0.88$ $0.83$ / $0.61$ $0.83$ / $0.32$ $0.83$ / $0.08$ $0.83$ / $0.08$
gamma+epsilon $0.83$ / $0.88$ $0.83$ / $0.88$ $0.83$ / $0.62$ $0.83$ / $0.32$ $0.83$ / $0.08$ $0.83$ / $0.08$
alpha1beta0 $0.83$ / $0.88$ $0.83$ / $0.88$ $0.83$ / $0.62$ $0.83$ / $0.32$ $0.83$ / $0.08$ $0.83$ / $0.08$
alpha2beta1 $0.83$ / $0.88$ $0.83$ / $0.88$ $0.83$ / $0.62$ $0.83$ / $0.32$ $0.83$ / $0.08$ $0.83$ / $0.08$
(g) Top 100 Neurons’ Relevance
$r$ $M_{p}$ $A_{g}$/$A_{t}$ $0\%$ $2\%$ $4\%$ $8\%$ $10\%$ $12\%$
epsilon $0.83$ / $0.88$ $0.83$ / $0.85$ $0.83$ / $0.56$ $0.83$ / $0.18$ $\mathbf{0.83}$ / $\mathbf{0.00}$ $\mathbf{0.83}$ / $\mathbf{0.00}$
gamma+epsilon $0.83$ / $0.88$ $0.83$ / $0.85$ $0.83$ / $0.56$ $0.83$ / $0.18$ $\mathbf{0.83}$ / $\mathbf{0.00}$ $\mathbf{0.83}$ / $\mathbf{0.00}$
alpha1beta0 $0.83$ / $0.88$ $0.83$ / $0.85$ $0.83$ / $0.56$ $0.83$ / $0.18$ $\mathbf{0.83}$ / $\mathbf{0.00}$ $\mathbf{0.83}$ / $\mathbf{0.00}$
alpha2beta1 $0.83$ / $0.88$ $0.83$ / $0.85$ $0.83$ / $0.56$ $0.83$ / $0.18$ $\mathbf{0.83}$ / $\mathbf{0.00}$ $\mathbf{0.83}$ / $\mathbf{0.00}$
(h) Top 150 Neurons’ Relevance
$r$ $M_{p}$ $A_{g}$/$A_{t}$ $0\%$ $2\%$ $4\%$ $8\%$ $10\%$ $12\%$
epsilon $0.83$ / $0.88$ $0.83$ / $0.88$ $0.83$ / $0.56$ $0.83$ / $0.18$ $\mathbf{0.83}$ / $\mathbf{0.00}$ $\mathbf{0.83}$ / $\mathbf{0.00}$
gamma+epsilon $0.83$ / $0.88$ $0.83$ / $0.88$ $0.83$ / $0.56$ $0.83$ / $0.18$ $\mathbf{0.83}$ / $\mathbf{0.00}$ $\mathbf{0.83}$ / $\mathbf{0.00}$
alpha1beta0 $0.83$ / $0.88$ $0.83$ / $0.88$ $0.83$ / $0.56$ $0.83$ / $0.18$ $\mathbf{0.83}$ / $\mathbf{0.00}$ $\mathbf{0.83}$ / $\mathbf{0.00}$
alpha2beta1 $0.83$ / $0.88$ $0.83$ / $0.88$ $0.83$ / $0.56$ $0.83$ / $0.18$ $\mathbf{0.83}$ / $\mathbf{0.00}$ $\mathbf{0.83}$ / $\mathbf{0.00}$
(i) Top 200 Neurons’ Relevance
V-D2 Different Amount of Perturbed Neurons
After analyzing the same number of neurons, perturbing different quantities of neurons can also impact unlearning performance.
Table 6(a) demonstrates the performance of our unlearning method on the MNIST dataset when analyzing the top 50 neurons’ relevance. Specifically, for $r=\text{epsilon}$, as the unlearning percentage $M_{p}$ increases, the accuracy $A_{t}$ decreases steadily. At $M_{p}=16\%$, the value of $A_{t}$ drops significantly to $0.005$, indicating that the unlearning process stabilizes and effectively erases the targeted information. Beyond this point, further increases in $M_{p}$ do not result in a significant change, as seen with $M_{p}=20\%$, where $A_{t}$ reaches $0.001$.
On the CIFAR-10 dataset, using the epsilon rule, $A_{t}$ (target class accuracy) shows a significant decrease as $M_{p}$ increases, while $A_{g}$ (non-target class accuracy) remains stable around $0.98$ or $0.97$. In Table 6(c), specifically, at $M_{p}=0\%$, $A_{t}=0.99$, indicating near-perfect accuracy for the target class. As $M_{p}$ increases to $10\%$ and $15\%$, $A_{t}$ drops sharply to $0.67$ and $0.34$. At $M_{p}=20\%$, $A_{t}$ further declines to $0.11$, and finally reaches its lowest values of $0.03$ and $0.06$ at $M_{p}=25\%$ and $30\%$, respectively. This trend demonstrates that as $M_{p}$ increases, the target class accuracy decreases progressively, achieving a strong unlearning effect.
Table 6(e) demonstrates the performance of applying our method to the CIFAR-100 dataset. Under the epsilon rule, $A_{t}$ decreases gradually as $M_{p}$ increases, while $A_{g}$ remains constant at $0.83$. Specifically, when $M_{p}$ increases from $0\%$ to $2\%$, $A_{t}$ slightly decreases from $0.88$ to $0.85$. As $M_{p}$ reaches $4\%$, $A_{t}$ drops more significantly to $0.56$, and further decreases to $0.18$ at $M_{p}=8\%$. Finally, at $M_{p}$ values of $10\%$ and $12\%$, $A_{t}$ reaches $0.00$, achieving complete unlearning.
A similar trend is observed in Table 6(h), which analyzes the top 150 neurons’ relevance on the mini-ImageNet dataset. For $r=\text{epsilon}$, the accuracy $A_{t}$ declines steadily as $M_{p}$ increases. When $M_{p}$ reaches $10\%$, $A_{t}$ drops to $0.00$, and even with further increases in $M_{p}$, this minimal accuracy remains, indicating that unlearning has effectively stabilized.
The trend is consistent across all relevance propagation rules $r$: as $M_{p}$ increases, $A_{t}$ decreases significantly, reaching a point where further unlearning does not substantially affect the results, indicating prosperous and stable unlearning.
Perturbing only a small subset of neurons fails to significantly affect a model’s memory because of the model’s redundancy. Information is often encoded across multiple neurons and layers in neural networks, especially those trained on complex tasks. This redundancy means that specific data points or features are not isolated to just a few neurons; their memory is distributed throughout the network.
Our experimental results reinforce this idea. For example, when analyzing the top $50$ neurons on the MNIST dataset or the top $150$ neurons on the mini-ImageNet dataset, significant memory disruption only occurs when a sufficient percentage of neurons are affected ($M_{p}\geq 16\%$ for MNIST and $M_{p}\geq 10\%$ for mini-ImageNet). This indicates that perturbing a small portion of neurons is insufficient to overcome the model’s redundancy, as the remaining unaffected neurons can compensate, preserving the model’s memory.
V-D3 Different Amount of Analyzed Neurons
Analyzing different quantities of neurons related to unlearning data can impact unlearning performance.
When comparing Table 6(a) and Table 6(b), we can observe that when analyzing the top $50$ neurons, unlearning is effectively achieved at $M_{p}$ = $16\%$ for most methods, with a further increase in $M_{p}$ to $20\%$ improving the unlearning effect. However, when analyzing the top 100 neurons, similar unlearning results are only reached when $M_{p}$ is $20\%$ or $24\%$.
Specifically, when analyzing the top $50$ neurons, at $M_{p}$ = $16\%$, utilizing propagation rules like epsilon, gamma+epsilon, alpha1beta0, alpha2beta1, and PatternAttribution $S(x){+-}$ achieve successful unlearning, with the target accuracy ($A_{t}$) dropping close to 00, while the global accuracy ($A_{g}$) remains high. In contrast, when analyzing the top $100$ neurons, only the PatternAttribution $S(x){+-}$ method achieves effective unlearning at $M_{p}$ = $16\%$, while other methods require $M_{p}$ to reach $20\%$ or more to achieve a similar drop in $A_{t}$, even though $A_{g}$ stays consistent.
When the number of neurons analyzed increases to $100$, while it includes more neurons related to the unlearning data, it also introduces more weak or unrelated neurons. These unrelated neurons can interfere with the method’s statistical steps, making it harder to identify the neurons relevant to the unlearning data accurately. Therefore, when we analyze more neurons, a higher $M_{p}$ is needed to achieve effective unlearning.
We also compare Table 6(c) and Table 6(d), which demonstrate the performance of unlearning on the CIFAR-10 dataset by analyzing the top 150 and 200 relevance. In Table 6(c), which analyzes the relevance of the top 150 neurons, we observe that the unlearning effect becomes prominent as the $M_{p}$ increases. At $M_{p}=25\%$, the $A_{t}$ drops sharply to $0.03$, and further to $0.06$ at $M_{p}=30\%$, while the $A_{g}$ remains stable at $0.97$. These bolded values indicate that the unlearning requirements are effectively met, as the target test accuracy is significantly reduced while the generalization accuracy stays consistent. This shows a clear unlearning effect when removing relevant neurons from the decision process at higher dropout probabilities.
In Table 6(d), where the top 200 neurons are analyzed, the unlearning effect becomes even more pronounced. At $M_{p}=25\%$, $A_{t}$ drops to $0.01$, and at $M_{p}=30\%$, the accuracy on the target test set reaches $0.00$, signifying near-complete unlearning. Meanwhile, the $A_{g}$ decreases slightly to $0.95$. Comparing the two tables, the unlearning effect intensifies as the number of neurons being analyzed increases. This suggests that analyzing more neurons allows for more thorough removal of target class information, leading to a stronger unlearning outcome.
In Table 6(e), which analyzes the relevance of the top 150 neurons on the CIFAR-100 dataset, the unlearning effect becomes evident as the $M_{p}$ increases. When $M_{p}=10\%$, the $A_{t}$ drops sharply to $0.01$, while the $A_{g}$ remains stable at $0.81$. This shows that the method achieves the desired unlearning at this point. At $M_{p}=12.5\%$, the target test accuracy reaches $0.00$, indicating complete unlearning of the target class, while $A_{g}$ still holds at $0.81$. These bolded values highlight that unlearning requirements are successfully met, as the target test accuracy significantly decreases with minimal impact on global accuracy.
Similarly, the results are consistent in Table 6(f), which considers the top 200 neurons. At $M_{p}=10\%$, $A_{t}$ drops to $0.01$, and at $M_{p}=12.5\%$, it reaches $0.00$, signifying a full unlearning effect. The generalization accuracy ($A_{g}$) remains at $0.81$ throughout, indicating that unlearning is effectively isolated to the target class without disrupting the overall performance. Comparing the two tables, increasing the number of neurons analyzed leads to a consistent unlearning trend, with unlearning performance stabilizing as more neurons are involved.
When comparing Table 6(g), Table 6(h), and Table 6(i), which examine the performance of unlearning on the mini-ImageNet dataset by analyzing the top 100, 150, and 200 relevance, we can observe some significant differences. The experimental results using other rules for $r$ show similar trends to those observed when $r$ is set to epsilon. Therefore, we will mainly focus on the experimental results and their changes when using epsilon as a propagation rule.
In Table 6(g), where the top $100$ neurons’ relevance is analyzed using $r=$ epsilon, the method begins to show an unlearning effect at $M_{p}=4\%$, with a noticeable drop in target accuracy $A_{t}$ to $0.61$. As $M_{p}$ increases to $8\%$, $A_{t}$ further decreases to $0.32$, and at $M_{p}=10\%$, it drops significantly to $0.08$. By $M_{p}=12\%$, $A_{t}$ stabilizes at $0.08$, while the global accuracy $A_{g}$ remains at $0.83$.
Building on this, in Table 6(h), the unlearning effect is more pronounced even at a lower $M_{p}$ when we analyze the top $150$ neurons with $r=$ epsilon. Here, $A_{t}$ drops to $0.56$ at $M_{p}=4\%$. As $M_{p}$ reaches $8\%$, $A_{t}$ decreases to $0.18$. At both $M_{p}=10\%$ and $M_{p}=12\%$, $A_{t}$ is reduced to 00, indicating complete unlearning, while $A_{g}$ remains consistent at $0.83$.
Similarly, in Table 6(i), where the top $200$ neurons’ relevance is analyzed with $r=$ epsilon, the trends are similar to those observed with the top $150$ neurons. Unlearning starts at $M_{p}=4\%$ with $A_{t}$ dropping to $0.56$, and as $M_{p}$ increases to $8\%$, $A_{t}$ falls to $0.18$. At both $M_{p}=10\%$ and $M_{p}=12\%$, $A_{t}$ reaches 00, while $A_{g}$ remains stable at $0.83$.
These results indicate that increasing the number of neurons analyzed from $100$ to $150$ leads to more effective unlearning at slightly lower $M_{p}$ values. However, increasing the number of neurons to $200$ does not enhance the unlearning effect beyond what is achieved with $150$ neurons. This pattern holds true across all relevance rules, demonstrating that the specific choice of $r$ has minimal impact on the method’s overall effectiveness.
V-E Experiment Analysis
V-E1 Impact of Neural Network Redundancy on the Number of Perturbed Neurons
Model parameter redundancy refers to the presence of a large number of parameters in deep learning models, where some parameters may not be essential for the model’s performance. This redundancy is especially notable in deep learning models, which are typically over-parameterized, meaning the number of parameters far exceeds the amount of training data needed. The redundancy also presents certain challenges when implementing machine unlearning.
When considering models trained on datasets with different numbers of classes but identical parameter counts, an interesting observation arises. Models trained on datasets with fewer classes, such as MNIST or CIFAR-10, tend to have higher redundancy. In these cases, it may require perturbing $16\%$ to $20\%$ of neurons in the fully connected layers to achieve the goal of unlearning.
Many parameters are redundant in models with high redundancy and contribute less to the model’s performance. Therefore, to forget specific information, it may be necessary to perturb a larger number of neurons to remove all relevant information. In other words, high redundancy increases the number of neurons that need to be perturbed, as individual perturbations may not significantly affect the model’s overall performance.
On the other hand, models trained on datasets with more classes, like CIFAR-100 or mini-ImageNet, exhibit lower redundancy. In such models, perturbing only $10\%$ of neurons in the fully connected layers is often sufficient for unlearning. This happens because each parameter becomes more critical to performance with more classes, making the model more sensitive to even small perturbations.
Each parameter contributes more significantly to the model’s performance in models with low redundancy. Since each parameter and neuron plays a more crucial role in such models, perturbing these parameters will directly impact the model’s behavior and memory. Consequently, in low redundancy models, perturbing a small number of neurons can substantially alter the model’s memory and achieve the goal of machine unlearning. This implies that low redundancy models may be more efficient in implementing machine unlearning because fewer perturbations are needed.
V-E2 Impact of Neural Network Redundancy on the Model Utility
Unlearning methods, particularly those targeting the perturbation of specific neuron weights within a neural network, have been shown to have minimal adverse effects on overall model utility. This phenomenon can be attributed to the inherent redundancy present within neural networks. Neural networks are designed with multiple layers and numerous neurons per layer, enabling them to distribute the task of data inference across a vast array of pathways. Consequently, the interference with a small subset of neuron weights does not singularly disrupt the network’s functionality.
The redundancy in neural networks means that no single pathway is solely responsible for the model’s ability to infer and classify data. Instead, the model’s predictive performance arises from the collective interactions of numerous neurons and pathways. This architectural characteristic ensures that the perturbation of a few specific neurons or parameters does not drastically affect the model’s capacity to generalize from the data. Such robustness is crucial, especially when considering the need for selective unlearning or model editing techniques to remove specific information or biases.
Additionally, neurons within a neural network exhibit differential sensitivity to various data types. The neurons activated by one input type may differ from those triggered by another. This differential activation pattern implies that altering the weights associated with specific neurons will not uniformly impact the model’s utility across all data types. For instance, modifying weights connected to neurons predominantly involved in recognizing features specific to one class of data may leave the recognition capabilities for other classes largely unaffected.
VI Conclusion
In this paper, we present a novel approach to machine unlearning via Layer-wise Relevance Analysis and Neuronal Path Perturbation. Layer-wise Relevance Analysis provides a detailed understanding of the internal workings of the neural network during the unlearning process. Assessing the relevance of each neuron to the unlearning information allows for more precise identification of the neurons that need to be perturbed. Neuronal Path Perturbation further enhances the unlearning process. By perturbing only the highly relevant neurons, it effectively disrupts the model’s memory of the specific data without overwriting or unnecessarily modifying a large portion of the network. Compared with traditional methods, our method addresses critical challenges of machine unlearning methods, such as lack of explanation, privacy guarantees, and the balance between unlearning effectiveness and model utility.
Experimental results on image classification models demonstrate the effectiveness of the proposed method in removing targeted data while maintaining overall model performance. Our method contributes to privacy-preserving machine learning and provides a practical solution for handling data deletion requests and safeguarding user privacy. Future work may focus on further optimizing the method and exploring its applications in more complex scenarios.